OpenStudy (anonymous):

Simplify: 1) (-64)^1/3 2) (25b^4)^1/2

jimthompson5910 (jim_thompson5910):

Hint: \[\LARGE \left(-64 \right )^{\frac{1}{3}} = \sqrt[3]{-64}\]

OpenStudy (anonymous):

-4*-4 = 16* -4 = -64

OpenStudy (anonymous):

so thats why its cubed -4^3 = -64